OpenStudy (anonymous):

. What is the product of 4/5 x 2 3/4 written in simplest form? * A. 1 1/5 B. 2 3/5 C. 1 2/3 D. 2 1/5

OpenStudy (anonymous):

D

OpenStudy (welshfella):

converting 2 3/4 to improper fraction gives 11/4 so we have 4/5 * 11/4 = 44/20 = 2 4/20 = 2 1/5

OpenStudy (anonymous):

Because 4/5 times 2 3/4 = 11/4 multiply across to get 44/20 simply and get 2 4/20 simply again and get 2 1/5

OpenStudy (welshfella):

I'm glad you gave that explanation

OpenStudy (anonymous):

thank you very much ,I have more